Sol Summary – March 21st

Crew 309 Sol Summary Report 21Mar2025

Sol: 5

Summary Title: Final Day

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • We conducted 3 EVAs today. GreenHab and Engineer went in the morning, commander A and Medic in the afternoon, and commander B and Scientist last. The crew began the cleaning process completing the RAM, GreenHab, and Food Inventory tasks this evening. Simulation ended after dinner and the crew was able to explore some of the Martian terrain outside of our EVA suits.

Look Ahead Plan: We will be packing up and heading home in the morning.

Anomalies in work:

  • During prep for our second EVA of the day, we determined that the helmet had become detached from the spacesuit worn by crew member one. Engineer, Medic, and Commander A were able to repair the suit for the remainder of the EVAs. A more permanent solution will be required for future missions.

Weather: Windy, 34F / 62F

Crew Physical Status: Nominal

EVA: 3 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report, Mission Summary

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site